# DyGMamba: Efficiently Modeling Long-Term Temporal Dependency on Continuous-Time Dynamic Graphs with State Space Models

### v3

@@ 5 Conclusion
We propose DyGMamba, an efficient CTDG representation learning model that can capture long-term temporal dependencies. DyGMamba first leverages a node-level SSM to encode long sequences of historical node interactions. It then employs a time-level SSM to learn edge-specific temporal patterns. The learned patterns are used to select the critical part of the encoded temporal information. DyGMamba achieves superior performance on dynamic link prediction, and moreover, it shows high efficiency and strong scalability compared with previous CTDG methods, implying a great potential in modeling huge amounts of temporal information with a limited computational budget. To supplement, we discuss the limitation of our work in App. K.
@@ G.1 Efficiency Statistics for all baselines
We provide the efficiency statistics for all baselines in Table 15.
@@
Datasets LastFM Enron MOOC UCI
Models # Params Time Mem # Params Time Mem # Params Time Mem # Params Time Mem
JODIE 0.75 4.4 2.28 0.75 0.07 1.30 0.75 0.78 2.36 0.75 0.03 1.44
DyRep 2.64 6.6 2.29 2.64 0.10 1.34 2.64 0.88 2.38 2.64 0.05 1.51
TGAT 4.02 22.75 4.15 4.02 1.28 3.46 4.02 4.08 3.64 4.02 0.60 3.42
TGN 3.68 12.14 2.21 3.68 0.15 1.45 3.68 1.03 2.54 3.68 0.08 1.51
CAWN 15.35 99.00 14.92 15.35 2.62 4.03 15.35 13.45 8.02 15.35 1.95 9.40
TCL 3.37 6.23 3.04 3.37 0.30 2.51 3.37 1.00 2.49 3.37 0.13 2.00
GraphMixer 2.45 16.35 2.78 2.45 1.20 2.23 2.45 4.02 2.40 2.45 0.73 2.19
DyGFormer 5.56 47.00 7.57 4.80 2.73 3.23 4.80 8.32 3.35 4.15 0.62 2.30
CTAN 0.45 3.33 1.44 0.47 0.50 1.33 0.68 3.22 2.30 0.50 0.38 1.30
DyGMamba 2.78 28.45 4.17 2.03 2.05 2.74 1.65 4.88 2.48 1.37 0.60 1.93
[…]
Datasets LastFM Enron MOOC UCI
[…]
DyGFormer 47.00 49.60 2331.20 2.73 32.80 89.54 8.32 64.20 534.14 0.62 34.80 21.58
CTAN 3.33 635.00 2114.55 0.50 173.00 86.50 3.22 138.00 444.36 0.38 236.00 89.68
DyGMamba 28.45 11.80 335.71 2.05 33.00 67.65 4.88 38.00 185.44 0.60 28.00 16.80
@@ Appendix I Motivation of Using SSM for Temporal Pattern Modeling
Apart from the analysis on Enron, we further analyze the impact of patch size on another long-term temporal dependent dataset MOOC. Fig. 7b shows the numbers of parameters of DyGFormer and DyGmamba with different patch sizes on MOOC. We confirm that patching greatly affects model sizes. We decrease the patch size gradually from 4 to 1 (the optimal value of DyGMamba’s hyperparameter $\rho\ \&\ p$ is 128& 41284128\ \&\ 4128 & 4) and track DyGMamba’s performance (Fig. 7a) as well as efficiency (Fig. 7b to 7d) on MOOC. Meanwhile, we also keep track on DyGFormer under the same patch size for comparison. Same as our analysis on Enron, we plot the performance of Variant A under different patch sizes in Fig. 7a as well. We can draw the same conclusions as in our analysis on Enron.
Appendix K Limitation
One limitation of our work is that DyGMamba is designed to only model CTDGs. DTDGs are represented as a sequence of graph snapshots, where all the edges in a snapshot are taken as existing simultaneously. This poses a challenge to DyGMamba because it can only encode edges sequentially, which are not suitable for modeling concurrent edges. A possible solution is to first employ methods such as graph neural networks (GNN) to encode each graph snapshot and then use SSM to model the temporal graph dynamics. However, the introduction of GNNs inevitably requires more computational resources, which would lower model efficiency.
Appendix L Performance of DyGMamba on Discrete-Time Dynamic Graphs

### v4

@@ 6 Conclusion
We propose DyGMamba, an efficient CTDG representation learning model that can capture long-term temporal dependencies. DyGMamba first leverages a node-level SSM to encode long sequences of historical node interactions. It then employs a time-level SSM to learn edge-specific temporal patterns. The learned patterns are used to select the critical part of the encoded temporal information. DyGMamba achieves superior performance on dynamic link prediction, and moreover, it shows high efficiency and strong scalability compared with previous CTDG methods, implying a great potential in modeling huge amounts of temporal information with a limited computational budget.
7 Limitation
One limitation of our work is that DyGMamba is designed to only model CTDGs. DTDGs are represented as a sequence of graph snapshots, where all the edges in a snapshot are taken as existing simultaneously. This poses a challenge to DyGMamba because it can only encode edges sequentially, which are not suitable for modeling concurrent edges. A possible solution is to first employ methods such as graph neural networks (GNN) to encode each graph snapshot and then use SSM to model the temporal graph dynamics (similar to approaches used in recent works on temporal knowledge graph modeling, e.g., Han et al. (2021)). However, the introduction of GNNs inevitably requires more computational resources, which would lower model efficiency.
Acknowledgements
Zifeng Ding receives funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation programme grant AVeriTeC (Grant agreement No. 865958). He also acknowledges travel support from the European Union’s Horizon 2020 research and innovation programme under ELISE Grant Agreement No 951847. Yuan He is supported by Samsung Research UK (SRUK), EPSRC projects UK FIRES (EP/S019111/1) and ConCur (EP/V050869/1). Michael Bronstein and Antonio Norelli are supported by EPSRC Turing AI World-Leading Research Fellowship No. EP/X040062/1, EPSRC AI Hub on Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI No. EP/Y028872/1. Antonio Norelli is also supported by Project CETI. Jingcheng Wu is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - SFB 1574 - Project number 471687386. This work is partially supported by the Munich Center for Machine Learning and is performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research. The authors also gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this project by providing computing time on the GCS Supercomputer HLRS at High-Performance Computing Center Stuttgart.
@@ G.1 Efficiency Statistics for all baselines
We provide the efficiency statistics for all baselines in Table 15. To supplement, in Figure 6, we further plot the comparison of model size, per epoch training time and GPU Memory across models on Reddit and Social Evo., analogous to the analysis in Sec. 4.3.1. CTAN cannot be trained on Social Evo. before timeout and therefore does not appear in the plot of Social Evo.. The complete statistics including the numbers in the plots can be found in Table 15.
@@
Datasets LastFM Enron MOOC UCI Reddit Social Evo.
Models # Params Time Mem # Params Time Mem # Params Time Mem # Params Time Mem # Params Time Mem # Params Time Mem
JODIE 0.75 4.4 2.28 0.75 0.07 1.30 0.75 0.78 2.36 0.75 0.03 1.44 0.75 3.95 1.10 0.75 4.70 1.71
DyRep 2.64 6.6 2.29 2.64 0.10 1.34 2.64 0.88 2.38 2.64 0.05 1.51 2.64 5.75 1.21 2.64 7.55 1.76
TGAT 4.02 22.75 4.15 4.02 1.28 3.46 4.02 4.08 3.64 4.02 0.60 3.42 4.02 16.33 2.98 4.02 25.50 3.89
TGN 3.68 12.14 2.21 3.68 0.15 1.45 3.68 1.03 2.54 3.68 0.08 1.51 3.67 2.05 1.67 3.67 3.83 1.78
CAWN 15.35 99.00 14.92 15.35 2.62 4.03 15.35 13.45 8.02 15.35 1.95 9.40 15.35 20.16 5.89 15.35 85.66 8.14
TCL 3.37 6.23 3.04 3.37 0.30 2.51 3.37 1.00 2.49 3.37 0.13 2.00 3.37 2.25 1.82 3.37 5.05 2.48
GraphMixer 2.45 16.35 2.78 2.45 1.20 2.23 2.45 4.02 2.40 2.45 0.73 2.19 2.44 4.92 1.57 2.45 15.50 2.71
DyGFormer 5.56 47.00 7.57 4.80 2.73 3.23 4.80 8.32 3.35 4.15 0.62 2.30 4.24 7.00 2.42 4.14 20.00 2.77
CTAN 0.45 3.33 1.44 0.47 0.50 1.33 0.68 3.22 2.30 0.50 0.38 1.30 0.53 0.86 1.54 0.45 2.41 0.63
DyGMamba 2.78 28.45 4.17 2.03 2.05 2.74 1.65 4.88 2.48 1.37 0.60 1.93 3.32 6.30 2.07 3.22 17.80 2.59
Figure 6: Efficiency comparison on Reddit and Social Evo. among DyGMamba and five baselines in terms of number (#) of parameters, training time per epoch and GPU memory.
@@ G.2 Total Training Time Comparison among DyGFormer, CTAN and DyGMamba
[…]
Datasets LastFM Enron MOOC UCI Reddit Social Evo.
[…]
DyGFormer 47.00 49.60 2331.20 2.73 32.80 89.54 8.32 64.20 534.14 0.62 34.80 21.58 4.24 24.60 104.30 20.00 30.22 604.40
CTAN 3.33 635.00 2114.55 0.50 173.00 86.50 3.22 138.00 444.36 0.38 236.00 89.68 0.53 327.18 173.41 0.45 Timeout Timeout
DyGMamba 28.45 11.80 335.71 2.05 33.00 67.65 4.88 38.00 185.44 0.60 28.00 16.80 3.32 26.80 88.98 17.80 24.40 434.32
@@ Appendix I Motivation of Using SSM for Temporal Pattern Modeling
Apart from the analysis on Enron, we further analyze the impact of patch size on another long-term temporal dependent dataset MOOC. Fig. 8b shows the numbers of parameters of DyGFormer and DyGmamba with different patch sizes on MOOC. We confirm that patching greatly affects model sizes. We decrease the patch size gradually from 4 to 1 (the optimal value of DyGMamba’s hyperparameter $\rho\ \&\ p$ is 128& 41284128\ \&\ 4128 & 4) and track DyGMamba’s performance (Fig. 8a) as well as efficiency (Fig. 8b to 8d) on MOOC. Meanwhile, we also keep track on DyGFormer under the same patch size for comparison. Same as our analysis on Enron, we plot the performance of Variant A under different patch sizes in Fig. 8a as well. We can draw the same conclusions as in our analysis on Enron.
Appendix K Performance of DyGMamba on Discrete-Time Dynamic Graphs
@@
Appendix L Detailed Discussion of Motivation
The goal of this work is to introduce a model capable of effective and efficient reasoning over CTDGs using extensive historical information. Our motivation is supported by the following considerations:
Abundant Historical Information in Real-World Datasets.
We quantify the availability of historical neighbors across various datasets for transductive and inductive link prediction tasks under random, historical, and inductive NSS settings, as shown in Table 22. Observations indicate that real-world CTDG datasets, such as LastFM and Enron, typically possess extensive one-hop historical information. Although it is feasible to perform predictions based solely on recent interactions, we posit that effectively leveraging longer historical contexts can further enhance model performance.
Limitations of Existing Models in Handling Extensive Historical Information.
Two recent state-of-the-art methods, DyGFormer and CTAN, explicitly designed to capture long-range historical dependencies, have been considered in our study. Our experimental results reveal distinct limitations: DyGFormer suffers from high computational complexity, while CTAN exhibits comparatively weaker predictive performance. Consequently, neither model achieves an optimal balance between effectiveness and computational efficiency. Our proposed model, DyGMamba, overcomes these limitations by demonstrating superior predictive capability and efficiency. Extensive experiments position DyGMamba as the top ranked method across multiple benchmark datasets.
Proven Performance Enhancement with Increased Historical Information.
In Appendix G.3, we have included an experiment showing different models’ performance with a varying number of temporal neighbors. From Figure 7, we observe that, for datasets with long-range temporal dependencies (such as Enron), an increase in sampled neighbors positively impacts model performance, particularly for DyGMamba and DyGFormer. We also find that for other baseline models, increasing the number of sampled neighbors does not consistently lead to performance improvements. This suggests that the key issue is not whether increasing sampled neighbors is always beneficial for dynamic graph modeling, but rather whether a model is robust enough to effectively process larger amounts of historical information. To further support this, we conducted additional experiments on Can. Parl [^12] Although Can. Parl is a DTDG, it requires huge number of sampled historical neighbors for optimal modeling. Therefore, it is considered here to prove the performance enhancement with increased historical information. using DyGMamba with an increasing number of sampled temporal neighbors (64, 128, 256, 512, 1024 and 2048). The results, presented in Table 23, confirm that DyGMamba consistently benefits from an increasing neighbor count until reaching the large number of 2048, further demonstrating its ability to leverage long-range temporal dependencies.
Temporal Pattern Modeling is Game Changer.
As mentioned in introduction, we wish to capture edge-specific temporal patterns for better CTDG modeling. The motivation can be well supported with the following experiment. We provide a comparison between DyGMamba, DyGFormer, and a modified version of DyGFormer where Transformer is replaced with Mamba (referred to as Variant G). Our results in Table 24 show a performance drop for Variant G across all datasets compared to DyGFormer. Additionally, Variant G performs significantly worse than DyGMamba, with particularly notable declines on the synthetic datasets S1, S2, and S3. We attribute this to two reasons:
  • Intrinsic Limitations of Mamba: Previous study (Waleffe et al., 2024) has shown that Mamba is generally less effective than Transformer in tasks requiring strong copying or in-context learning abilities. In CTDG modeling, strong copying ability is crucial, as many predictions are based on recalling repeated edges. Consequently, replacing Transformer with Mamba in DyGFormer inevitably leads to a performance drop, despite the gains in efficiency.
  • The Importance of Temporal Pattern Modeling: Our dynamic information selection module plays a crucial role in enabling a Mamba-based model to outperform a Transformer-based model in CTDG modeling, especially when clear temporal patterns exist. This suggests that effectively capturing temporal patterns can significantly enhance the competitiveness of Mamba-based models in this context, further validating the novelty and the contribution of our work.
To summarize, the design our DyGMamba achieves a good balance between efficiency and effectiveness, thanks to temporal pattern modeling.
Table 22: Number of available historical neighbors for all datasets on both transductive and inductive link prediction under random/historical/inductive NSS settings. Trans LP and Ind LP denote transductive and inductive link prediction, respectively.
NSS Random NSS Historical NSS Inductive NSS
Trans LP Ind LP Trans LP Trans LP Ind LP
Datasets Avg. Max Avg. Max Avg. Max Avg. Max Avg. Max
LastFM 2,253.03 51,767 2,333.75 51,767 2,393.71 51,767 2,237.29 51,767 2,309.49 51,767
Enron 1,681.80 21,512 1,693.40 21,511 1,734.19 21,512 1,675.21 21,512 1,670.65 21,511
MOOC 2,304.04 19,473 2,312.62 19,473 3,265.62 19,473 2,293.64 19,473 2,275.23 19,473
Reddit 4,129.77 58,726 4,196.20 58,726 4,604.99 58,726 4,128.88 58,726 4,186.38 58,726
Wikipedia 144.39 1,937 149.67 1,937 164.49 1,937 144.17 1,937 149.22 1,937
UCI 172.82 1,546 193.87 1,546 261.31 1,546 173.08 1,546 192.72 1,546
Social Evo. 63,962.35 124,565 68,701.32 124,565 60,403.62 124,565 60,395.83 124,565 65,191.24 124,565
Table 23: DyGMamba’s performance with increasing sampled neighbors on transductive link prediction under random NSS.
Can. Parl 64 128 256 512 1024 2048
DyGMamba 72.58 74.46 76.85 82.67 94.88 99.57
Table 24: Comparison among Variant G, DyGFormer and DyGMamba on transductive link prediction under random NSS.
Models LastFM Enron MOOC Reddit Wikipedia UCI Social Evo. S1 S2 S3
Variant G 92.69 92.24 87.13 98.89 98.70 95.32 94.29 54.40 55.13 72.53
DyGFormer 92.95 92.42 87.66 99.22 99.03 95.74 94.63 55.19 57.80 79.20
DyGMmaba 93.35 92.65 89.21 99.32 99.15 95.91 94.77 81.85 85.36 86.59